OpenStudy (anonymous):

14(x-4)-(x+2)=(x+2)(x-4) When I work thhis out I come out with x^2 + 15x -46 though they showed me x^215x+50=0 What am I doing wrong?

OpenStudy (anonymous):

14x-56-x-2=x^2-2x-8 , 13x-58=x^2-2x-8 x^2-15x+50=0

OpenStudy (anonymous):

jithu 1984 Thank you , but I have a question when you bring the x^2 -2x -8 to the left side don't you have to cancel out on the right side to move to the left by changing them to the oppisite like + to a - or am I wrong

OpenStudy (anonymous):

moving +13x to Right changes it to -13x ,which adds-up to give -15x and -58 changes to +58 when moved to other side of equal sign, adds to-8 and give +50.

OpenStudy (anonymous):

Thsnk you, this is what gets me i am always second guessing everything again thank you.